Two light, conducting spheres, each 6 mm diameter and having a mass of 10 mg, are

suspended from the same point by fine insulated fibers of 50cm long. Due to electrostatic

repulsion the spheres are in equilibrium when 3 cm apart.


1
Expert's answer
2021-05-07T09:45:09-0400

"sin\Big(\dfrac{\theta}{2}\Big)=\dfrac{1.5}{50}"

"\dfrac{\theta}{2}=sin^{-1}\Big(\dfrac{1.5}{50}\Big)"

"\theta=3.43\degree"


Balancing forces in x direction,

"Tcos(90\degree-1.715\degree)=F"

"T=\dfrac{F}{cos88.28\degree}\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(1)"


Balancing forces in y direction

"Tsin88.28\degree=mg"

"T=\dfrac{mg}{sin88.28\degree}\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space\space(2)"


From equation (1) and (2)

"\dfrac{F}{mg}=cot88.28\degree"

"F=mg\times cot88.28\degree"

"F=2.94\times10^{-3}\space N"


Let charges on both the sphere be "q"

Electrostatic force, "F=\dfrac{kq^2}{r^2}"

"q=\sqrt{\dfrac{(3\times10^{-2})^2(2.94\times10^{-3})}{9\times10^{9}}}"

"q=1.714\times10^{-8}\space C"